OpenStudy (anonymous):

Please help! 1.For which value of x does the graph of y = 3x2 − 2x − 5 cross the x-axis? 2.What are the approximate solutions of 2x2 + 9x = 8 to the nearest hundredth?

OpenStudy (campbell_st):

an easy method for both problems is to use the general quadratic formula

OpenStudy (anonymous):

i dont know how... i was sick for the week they tought this

OpenStudy (campbell_st):

ok... for a quadratic \[ax^2 + bx + c = 0\] the roots of zeros can be found using \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (campbell_st):

in your 1st question a = 3, b = -2 and c = -5 substitute them and then calculate the 2 answers

OpenStudy (anonymous):

so... x=-1/3?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

not quite when you substitute its \[x = \frac{2 \pm \sqrt{(-2)^2 - 4 \times 3 \times (-5)}}{2 \times 3} = \frac{2 \pm 8}{6}\] so you have 2 solutions \[x = \frac {2 + 8}{6}~~~~~and~~~~~x = \frac{2 - 8}{6}\] just simplify them

OpenStudy (anonymous):

so.. -1?

OpenStudy (anonymous):

-1 one is correct right @campbell_st

OpenStudy (campbell_st):

that's correct... and the other is..?

OpenStudy (anonymous):

i cant figure it out...

OpenStudy (anonymous):

2^{2} +9x =8

OpenStudy (anonymous):

can you please show me how to do it @campbell_st ?

OpenStudy (campbell_st):

add you needed to do for the 2nd intercept is simplify \[x = \frac{-2 -8}{6} = \frac{-10}{6} ~~~so~~~~x = \frac{-5}{3} \]

OpenStudy (campbell_st):

(2) rewrite the equation \[2x^2+ 9x - 8 = 0\] so using the formula you get \[x = \frac{-9 \pm \sqrt{9^2 - 4 \times 2 \times -8}}{2 \times 2}\] so there are 2 solutions \[x = \frac{-9 + \sqrt{9^2 - 4 \times 2 \times -8}}{2 \times2}~~and~~~x = \frac{-9 - \sqrt{9^2 - 4 \times 2 \times -8}}{2 \times 2}\] just calculate the answers

OpenStudy (campbell_st):

the queston says the solution to the nearest hundredth... so use a calculator and the round the answers to 2 decimal places.

OpenStudy (campbell_st):

so you need to solve to 2 decimal places \[x = \frac{-9 + \sqrt{145}}{4}~~~and~~~ x = \frac{-9 - \sqrt{145}}{4}\]